substitution of y=-2x+3 3x+y=2
Answer by Menjax(62)

y=-2x+3 3x+y=2
sub y=-2x+3 into 3x+y=2:
3x + (-2x+3) = 2
X + 3 = 2
x = -1
sub x =-1 into y=-2x+3
y=-2*-1+3
y = 5
so x is -1 and y is 5